Fermi National Laboratory
 

Operation Summary for August 30, 2010 - September 1, 2010

 

Monday August 30

The day shift began with the Tevatron (TeV) orbit checks, with Operations monitoring the Recycler stash (211.5E10), with the Antiproton Source (Pbar) stacking (15.6E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, TeV orbit checks continued from the previous shift. Operations put the TeV into HEP shot setup at 9:30 AM and established store 8042 at 10:45 AM with an average initial luminosity (AIL) of 215E30. Pbar personnel repaired a Debuncher RF station (DRF1-7) with only a brief interruption to stacking. EE Support personnel swapped out the regulator cards for two MI quadrupole power supplies (QC330 and QT402) and installed diagnostics on F17B3 to investigate regulation problems.

During the evening shift, Operations terminated store 8042 at 6 PM. Operators prepared the CDF and D-Zero collision halls for accesses. Operations put the TeV into shot setup at 9:45 PM and established store 8045 at 11 PM with an AIL of 194E30. A ventilation issue at AP0 inhibited beam to the Antiproton Source for almost two hours. ES&H and Mechanical Support personnel investigated and found that an exhaust fan had tripped off.

Tuesday August 31

The midnight shift began with Operations monitoring store 8045 (70.72E30) and the stash (231.6E10), with Pbar stacking (11.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations delivered beam to the Pbar, NuMI, MiniBooNE, and, when available, transferred antiprotons to the Recycler.

During the day shift, Operations terminated store 8045 at noon, put the TeV into shot setup, and established store 8046 at 1:30 PM with an AIL of 236E30. Instrumentation personnel calibrated the Recycler DCCT after the antiprotons were sent to the TeV.

During the evening shift, Operations reported that the NuMI extraction kicker faulted at a high voltage output cable termination. Operators called in EE Support personnel to reterminate the cable; Operations reestablished the beam to NuMI at 9:15 PM. The MI-52 area developed a vacuum problem that recovered very slowly.

Wednesday Sept. 1

The midnight shift began with Operations monitoring store 8046 (75.24E30) and the stash (217.1E10), with Pbar stacking (16.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the MI-520 vacuum problems continued; ion pumps tripped off multiple times. Operations terminated store 8046 at 6 AM and put the TeV into shot setup, but this was hindered by the MI vacuum problems. Operations halted stacking and beam to NuMI during shot setup to help keep the vacuum stable and increase the chances of injecting the antiprotons into the TeV.

The Plans for Wednesday

The plans for today are to monitor the store. The Run Coordinator will allow MI personnel and vacuum technicians to access into MI to find and repair the vacuum problem. She will also allow other machines to access for maintenance work.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/01/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory